2) Find the maximum revenue if the price of selling x hammers is p(x) = 2 – 3x
and independent. 14) Graph: - x + 12 = 1 +
Answer #1

Plac)=2-30 Revenue (Rears) = plac). R (a) = (2-30) 2 2x - 3x To maximize revenue, set R (a)=0 2 - 6x = 0 6x=2 3 2. R ( 33 ) 2E 10 (0,3) -10 0 10 (0, -3) -10